OpenStudy (anonymous):

If the ratio of n/m = 7/8 and the ratio of m/p = 4/5, what is the ratio of n/p?

OpenStudy (phi):

do you know that \[ \frac{n}{m}\cdot\frac{m}{p}=\frac{n}{\cancel{m}}\cdot\frac{\cancel{m}}{p} = \frac{n}{p}\]

OpenStudy (anonymous):

oh wow didn't even notice. so i would just multiply 7/8 and 4/5 right?

OpenStudy (phi):

yes. what do you get?

OpenStudy (anonymous):

7/10

OpenStudy (phi):

yes. sometime you have to divide (flip one of the ratios), but these problems are not hard once you see the trick.

OpenStudy (anonymous):

k thanks :)